# Factoriality of groupoid von Neumann algebras

### v2

@@ 5.3 Kleppner’s condition and Factoriality
Suppose first that $E\subset\operatorname{Iso}(\mathcal{G})$ is central Borel set with $\mu_{s}(E)<\infty$ which is not contained in $\mathcal{G}^{(0)}$. Let $f:E\to\mathbb{C}$ be the function satisfying (17). Then a calculation shows that $f\cdot 1_{E}$ defines an element in $\mathcal{Z}(L_{\omega}(\mathcal{G}))$, which is moreover not contained in $L^{\infty}(\mathcal{G}^{(0)},\mu)$.
@@
Let $(\mathcal{G},\mu,\omega)$ be a twisted measured discrete groupoid such that $L_{\omega}(\mathcal{G})$ is a factor. Then $\mathcal{G}$ satisfies Kleppner’s condition.
@@
Suppose that $\mathcal{G}$ does not satisfy Kleppner’s condition. Let $A$ be an $\omega$-regular bisection with $\mu_{s}(\Omega_{A})<\infty$ and $\Omega_{A}\not\subset\mathcal{G}^{(0)}$. Let $\mathcal{A}:=\{(g,a)\in A\times\mathcal{G}\ |\ s(g)=t(a)\}$ and define
$f_{0}:\mathcal{A}\to\mathbb{T}:f(g,a)=\omega(g,a)\omega(ga,g^{-1}).$
@@
$\omega(g,x)\omega(gx,g^{-1})\overline{\omega(h,y)}\overline{\omega(hy,h^{-1})}$ $=\omega(g,x)\omega(gx,g^{-1})\overline{\omega(h,y)}\omega(gxg^{-1},h)$
$=\omega(g,x)\omega(gx,g^{-1}h)\omega(g^{-1},h)\overline{\omega(h,y)}$
$=\omega(g,xg^{-1}h)\omega(x,g^{-1}h)\omega(g^{-1},h)\overline{\omega(h,y)}$
$=\omega(g,g^{-1}hy)\omega(g^{-1}h,y)\omega(g^{-1}h)\overline{\omega(h,y)}$
@@
The map $(g,a)\mapsto gag^{-1}$ is a countable-to-one Borel surjection from $\mathcal{A}$ onto $\Omega_{A}$. Therefore there exists a Borel section $\sigma:\Omega_{A}\to\mathcal{A}$. We get then get that the formula $f(x)=f_{0}(\sigma(x))$ defines a Borel map which is independent of the choice of section. Additionally, another direct computation below shows that $f$ satisfies
@@
$\omega(hg,a)$ $\omega(hga,g^{-1}h^{-1})\overline{\omega(g,a)}\ \overline{\omega(ga,g^{-1})}\ \overline{\omega(h,gag^{-1})}\ \overline{\omega(hgag^{-1},h^{-1})}$
$=\omega(hg,a)\omega(hga,g^{-1})\overline{\omega(g^{-1},h^{-1})}\ \overline{\omega(g,a)}\ \overline{\omega(ga,g^{-1})}\ \overline{\omega(h,gag^{-1})}$
$=\omega(hg,ag^{-1})\omega(a,g^{-1})\overline{\omega(g^{-1},h^{-1})}\ \overline{\omega(g,a)}\ \overline{\omega(ga,g^{-1})}\ \overline{\omega(h,gag^{-1})}$
$=\omega(h,gag^{-1})\omega(g,ag^{-1})\omega(a,g^{-1})\overline{\omega(g,a)}\ \overline{\omega(ga,g^{-1})}\ \overline{\omega(h,gag^{-1})}$
@@
By construction $f\lambda_{\Omega_{A}}\in L_{\omega}(\mathcal{G})\setminus L^{\infty}(\mathcal{G}^{(0)},\mu)$. We claim that $f\lambda_{\Omega_{A}}\in\mathcal{Z}(L_{\omega}(\mathcal{G}))$. Let $B\subset\mathcal{G}$ be any Borel bisection. Since $B\cdot\Omega_{A}=\Omega_{A}\cdot B$, for any point $bx=yb\in B\cdot\Omega_{A}$ with $x,y\in\Omega_{A}$ we have
@@ 6.1 Group bundles
First assume that $\mu(Y)>0$. By Lemma 6.2 below there exists a section $\sigma\colon\mathcal{G}^{(0)}\rightarrow\mathcal{G}$ such that for every $x\in Y$ we have that $\sigma(x)\neq e_{x}$ and $\sigma(x)$ has finite conjugacy class. Then $\sigma(Y)\subset\mathcal{G}\setminus\mathcal{G}^{(0)}$ is a nontrivial bisection and $|C_{g}|<+\infty$ for every $g\in\sigma(Y)$ by construction.
@@
Let $(\mathcal{G},\mu)$ be a discrete Borel groupoid. Write $\Gamma_{x}=\{g\in\mathcal{G}:s(g)=t(g)\}$ for $x\in\mathcal{G}^{(0)}$ and for an element $g\in\Gamma_{x}$ write $C_{g}\subset\Gamma_{x}$ for its conjugacy class. The sets $B=\{x\in\mathcal{G}^{(0)}\mid\Gamma_{x}\text{ is trivial}\}$ and $A=\{x\in\mathcal{G}^{(0)}\mid\Gamma_{x}\text{ is icc}\}$ are Borel and there exists a Borel section $\sigma\colon\mathcal{G}^{(0)}\rightarrow\operatorname{Iso}(\mathcal{G})$ such that $\sigma(x)$ is nontrivial and has finite conjugacy class for every $x\in A\setminus B$.
@@
The source map $s\colon\operatorname{Iso}(\mathcal{G})\rightarrow\mathcal{G}^{(0)}$ is a surjective and countable-to-one Borel map. So by the Lusin–Novikov theorem there exists a countable collection of Borel subsets $U_{n}\subset\operatorname{Iso}(\mathcal{G})$ such that $\operatorname{Iso}(\mathcal{G})=\bigcup_{n}U_{n}$ and such that $s\colon U_{n}\rightarrow\mathcal{G}^{(0)}$ is a Borel isomorphism for every $n$. Then we have that $B=\bigcap_{n}(\mathcal{G}^{(0)}\cap U_{n})$, which is Borel.
The restriction of $s$ to $\operatorname{Iso}(\mathcal{G})\setminus\mathcal{G}^{(0)}$ is again countable-to-one, so we find countably many Borel section $\tau_{n}\colon X\setminus B\rightarrow\operatorname{Iso}(\mathcal{G})$ such that $\operatorname{Iso}(\mathcal{G})\setminus\mathcal{G}^{(0)}=\bigcup_{n}\tau_{n}(\mathcal{G}^{(0)})$. For $g\in\Gamma_{x}$ write $C_{g}\subset\Gamma_{x}$ for its conjugacy class. If $\sigma\colon\mathcal{G}^{(0)}\rightarrow\operatorname{Iso}(\mathcal{G})$ is a Borel section, then the subset $\bigsqcup_{x\in\mathcal{G}^{(0)}}C_{\sigma(x)}=\bigcup_{n}U_{n}\sigma(\mathcal{G}^{(0)})U_{n}^{-1}\subset\operatorname{Iso}(\mathcal{G})$ is Borel. Then, since for any countable-to-one Borel map $q$, the assignment $x\mapsto|q^{-1}(x)|$ is Borel, we get that $x\mapsto|C_{\tau_{n}(x)}|$ is Borel for every $n$. Now define Borel maps
$f_{n}\colon\mathcal{G}^{(0)}\setminus B\rightarrow\mathbb{N}\cup\{+\infty\}:\quad f(x)=\begin{cases}+\infty&\text{ if }|C_{\tau_{n}(x)}|=+\infty,\\ n&\text{ if }|C_{\tau_{n}(x)}|<+\infty,\end{cases}$ (18)
@@
$\sigma(x)=\begin{cases}\tau_{n}(x)&\text{ if }f(x)=n<+\infty,\\ e_{x}&\text{ if }x\notin A\setminus B.\end{cases}$ (19)
Then $\sigma(x)$ is nontrivial and has finite conjugacy class for every $x\in A\setminus B$. ∎
@@ 6.2 Transformation groupoids
As mentioned in the introduction Vaes provided necessary and sufficient conditions for when the crossed $L^{\infty}(X)\rtimes G$ by a nonsingular group action $G\curvearrowright(X,\mu)$ is a factor in [Vae20]. In this subsection we deduce this characterization as a corollary of Theorem A, see Corollary 6.4 below. In particular, [Vae20, Corollary] says that if the action $G\curvearrowright(X,\mu)$ is moreover pmp, the crossed product $L^{\infty}(X)\rtimes G$ is a factor if and only if $G\curvearrowright(X,\mu)$ is ergodic and every $g\in G$ satisfying $\mu(\{x\in X:g\cdot x=x\})>0$ has an infinite conjugacy class in $G$.
Recall that to a nonsingular action $G\curvearrowright(X,\mu)$ of a countable group $G$ on a standard probability space $(X,\mu)$ we associate the transformation groupoid $G\ltimes X$. As a set, we have that $G\ltimes X=G\times X$. The set of composable pairs is given by $(G\ltimes X)^{(2)}=\{((g,h\cdot x),(h,x))\mid g,h\in G,x\in X\}$ and multiplication is defined by $(g,h\cdot x)(h,x)=(gh,x)$. Identifying the unit space $(G\ltimes X)^{(0)}$ with $X$, we get that $s(g,x)=x$ and $t(g,x)=g\cdot x$. There is a canonical isomorphism $L^{\infty}(X)\rtimes G\cong L(G\ltimes X)$.
@@
First assume that $\mathcal{G}$ is not icc. Then there exists a nontrivial bisection $B\subset\operatorname{Iso}(\mathcal{G})\setminus\mathcal{G}^{(0)}$ which has $\mu$-finite conjugacy class $\Omega_{B}$. Define the function $F\colon X\rightarrow\mathcal{P}(G)$ by
@@
[…]
@@
Using that $L(G\ltimes X)\cong L^{\infty}(X)\rtimes G$ and that $L^{\infty}(X)^{G}\subset\mathcal{Z}(L^{\infty}(X)\rtimes G)$ we thus obtain the following as a Corollary of Theorem A and Proposition 6.3.
@@
    The crossed product $L^{\infty}(X)\rtimes G$ is a factor.
@@ 6.3 Full subgroupoids and partial actions
In the previous section we’ve demonstrated that factoriality of crossed products $L^{\infty}(X)\rtimes G$ is inherently linked to factoriality of certain subgroupoids of $G\ltimes X$. First we describe subgroupoids which share the same ideal structure as our original groupoid.
@@
For what follows, we let $\mathcal{R}$ denote the full equivalence relation on the set $\mathbb{N}$ of positive integers. Our goal is to show that subgroupoids induced from full subsets of $\mathcal{G}^{(0)}$ share the same ideal structure as $\mathcal{G}$. What follows is essentially the same proof as in [CRS17, Lemmas 2.2, 2.3, 2.4, Proposition 2.5] in the ample groupoid case.
@@
The proof is a back and forth construction using Lemma 6.6, and is identical to the proof in [CRS17, Lemma 2.4], so we omit the details here. ∎
@@
$\mathcal{G}\to\mathcal{G}|_{K}\times\mathcal{R}:g\mapsto W^{-1}gW$
gives us the groupoid isomorphism, with inverse $g\mapsto WgW^{-1}$. ∎
@@
This follows from the identity $L(\mathcal{G})\overline{\otimes}B(\ell^{2}(\mathbb{N}))=L(\mathcal{G}\times\mathcal{R})\cong L(\mathcal{G}|_{K}\times\mathcal{R})=L(\mathcal{G}|_{K})\overline{\otimes}B(\ell^{2}(\mathbb{N}))$. ∎
@@
If $G\curvearrowright X$ is induced from $H\curvearrowright Y$, then $Y$ is full by definition. Conversely, note that by fullness of $Y$, that $\bigcup_{g\in G}gY=t((G\ltimes X)Y)=X$. To show these sets are mutually disjoint, let $g\not\in H$. If there is some $y\in gY\cap Y$, then $(g,y)\in(G\ltimes X)|_{Y}=H\ltimes Y$. Thus, $g\in H$, reaching a contradiction. The measurable setting is the same, with equality replaced with equality up to a negligable subset. ∎
Note that subgroupoids of $G\ltimes X$ are not necessarily crossed products.
@@
The fix for this is to look at partial actions instead. The following is a translation of the partial dynamical system construction in the topological case to the Borel setting. See the book [Exe17] and the references therein for more details.
@@
Set $Y_{g}=\{x\in X_{g}\;|\;x\in Y,\sigma_{g^{-1}}(x)\in Y\}$. For any $x\in Y_{g}$, $\sigma_{g^{-1}}(x)\in Y_{g^{-1}}$. It follows that $\sigma_{g}\colon Y_{g}\rightarrow Y_{g^{-1}}$ is a bijection with inverse $\sigma_{g^{-1}}$. The only non-trivial fact to verify is that $(\sigma_{g}\circ\sigma_{h})(x)\in Y_{gh}$ if $x\in\sigma_{h}^{-1}(Y_{h}\cap Y_{g^{-}1})$. Since $(\sigma_{g}\circ\sigma_{h})(x)=\sigma_{gh}(x)\in Y\cap X_{gh}$, it is enough to verify that $x\in Y_{(gh)^{-1}}$. This follows from the fact that $x,\sigma_{gh}(x)\in Y$. Finally, to see that $(G\ltimes X)_{Y}=G\ltimes Y$, observe that $\{g\}\times Y_{g^{-1}}$ form a basis for both groupoids. ∎
Indeed, the partial actions always come from the restriction of a global action. The proof is a modification of [Aba03, Theorem 2.5] in the topological case.
@@
Let $G\curvearrowright_{\sigma}Y$ be a partial action on a standard Borel space $Y$. There is an embedding $Y\hookrightarrow X$ into a standard Borel space $X$ and a global action $G\curvearrowright_{\theta}X$ such that $G\ltimes Y=(G\ltimes X)|_{Y}$. Furthermore, $X$ can be chosen such that $Y$ is $G\ltimes X$-full.
@@
Next, we show by induction that $[(g,x)]\cap A$ is nonempty for every $(g,x)\in G\times X$. Take an arbitrary element $(g_{i},x)\in G\times X$. If $i=0$ we are done. Now assume that $i>0$. If $x\notin\cup_{k=0}^{i-1}Y_{g_{i}^{-1}g_{k}}$, then we have that $(g_{i},x)\in A$ by definition. If $x\in\cup_{k=0}^{i-1}Y_{g_{i}^{-1}g_{k}}$, there exists a $k<i$ such that $x\in Y_{g_{i}^{-1}g_{k}}$ and we get that $(g_{i},x)R(g_{k},\sigma_{g_{i}^{-1}g_{k}}(x))$. By the induction hypothesis we get that $[(g_{i},x)]\cap A=[(g_{k},\sigma_{g_{i}^{-1}g_{k}}(x))]\cap A\neq\emptyset$.

### v3

@@ 5.3 Kleppner’s condition and Factoriality
Suppose first that $E\subset\operatorname{Iso}(\mathcal{G})$ is central Borel set with $\mu_{s}(E)<\infty$ which is not contained in $\mathcal{G}^{(0)}$. Let $f:E\to\mathbb{C}$ be a function satisfying (17). Then a calculation shows that $f\cdot 1_{E}$ defines an element in $\mathcal{Z}(L_{\omega}(\mathcal{G}))$, which is moreover not contained in $L^{\infty}(\mathcal{G}^{(0)},\mu)$.
@@
Let $(\mathcal{G},\mu,\omega)$ be a twisted measured discrete groupoid. If $L_{\omega}(\mathcal{G})$ is a factor, then $\mathcal{G}$ satisfies Kleppner’s condition.
@@
Suppose that $\mathcal{G}$ does not satisfy Kleppner’s condition. Let $A$ be an $\omega$-regular bisection with $\mu_{s}(\Omega_{A})<\infty$ and $A\not\subset\mathcal{G}^{(0)}$. Note that the latter condition is equivalent to $\Omega_{A}\not\subset\mathcal{G}^{(0)}$. Let $\mathcal{A}:=\{(g,a)\in\mathcal{G}\times A\ |\ s(g)=t(a)\}$ and define
$f_{0}:\mathcal{A}\to\mathbb{T}:f_{0}(g,a)=\omega(g,a)\omega(ga,g^{-1}).$
@@
$\omega(g,x)\omega(gx,g^{-1})\overline{\omega(h,y)}\ {\color[rgb]{1,0,0}\overline{\omega(hy,h^{-1})}}$ $=\omega(g,x){\color[rgb]{1,0,0}\omega(gx,g^{-1})}\overline{\omega(h,y)}{\color[rgb]{1,0,0}\omega(gxg^{-1},h)}$
$={\color[rgb]{1,0,0}\omega(g,x)\omega(gx,g^{-1}h)}\omega(g^{-1},h)\overline{\omega(h,y)}$
$={\color[rgb]{1,0,0}\omega(g,xg^{-1}h)\omega(x,g^{-1}h)}\omega(g^{-1},h)\overline{\omega(h,y)}$
$={\color[rgb]{1,0,0}\omega(g,g^{-1}hy)\omega(g^{-1}h,y)}\omega(g^{-1},h)\overline{\omega(h,y)}$
@@
The map $(g,a)\mapsto gag^{-1}$ is a countable-to-one Borel surjection from $\mathcal{A}$ onto $\Omega_{A}$. Therefore there exists a Borel section $\sigma:\Omega_{A}\to\mathcal{A}$. We then get that the formula $f(x)=f_{0}(\sigma(x))$ defines a Borel map which is independent of the choice of section. Additionally, another direct computation below shows that $f$ satisfies
@@
$\omega(hg,a)$ ${\color[rgb]{1,0,0}\omega(hga,g^{-1}h^{-1})}\overline{\omega(g,a)}\ \overline{\omega(ga,g^{-1})}\ \overline{\omega(h,gag^{-1})}\ {\color[rgb]{1,0,0}\overline{\omega(hgag^{-1},h^{-1})}}$
$={\color[rgb]{1,0,0}\omega(hg,a)\omega(hga,g^{-1})}\overline{\omega(g^{-1},h^{-1})}\ \overline{\omega(g,a)}\ \overline{\omega(ga,g^{-1})}\ \overline{\omega(h,gag^{-1})}$
$={\color[rgb]{1,0,0}\omega(hg,ag^{-1})}\omega(a,g^{-1}){\color[rgb]{1,0,0}\overline{\omega(g^{-1},h^{-1})}}\ \overline{\omega(g,a)}\ \overline{\omega(ga,g^{-1})}\ \overline{\omega(h,gag^{-1})}$
$={\color[rgb]{1,0,0}\omega(h,gag^{-1})}\omega(g,ag^{-1})\omega(a,g^{-1})\overline{\omega(g,a)}\ \overline{\omega(ga,g^{-1})}\ {\color[rgb]{1,0,0}\overline{\omega(h,gag^{-1})}}$
@@
By construction $f\lambda_{\Omega_{A}}\in L_{\omega}(\mathcal{G})\setminus L^{\infty}(\mathcal{G}^{(0)},\mu)$. We now prove that $f\lambda_{\Omega_{A}}\in\mathcal{Z}(L_{\omega}(\mathcal{G}))$. Let $B\subset\mathcal{G}$ be any Borel bisection. For any $b\in B$ and $gag^{-1}\in\Omega_{A}$, such that $t(g)=s(b)$, we have
$b(gag^{-1})=bgag^{-1}s(b)=(bg)a(bg)^{-1}b\in\Omega_{A}\cdot B\;.$
Therefore $B\cdot\Omega_{A}\subseteq\Omega_{A}\cdot B$. Similarly, $\Omega_{A}\cdot B\subseteq B\cdot\Omega_{A}$ and thus $B\cdot\Omega_{A}=\Omega_{A}\cdot B$. For any element $bx=yb\in B\cdot\Omega_{A}$ with $x,y\in\Omega_{A}$ we have
@@ 6.1 Group bundles
First assume that $\mu(Y)>0$. Denote by $e_{x}$ the neutral element in $\Gamma_{x}$ for every $x$ in $X$. By Lemma 6.2 below there exists a section $\sigma\colon\mathcal{G}^{(0)}\rightarrow\mathcal{G}$ such that for every $x\in Y$ we have that $\sigma(x)\neq e_{x}$ and $\sigma(x)$ has finite conjugacy class. Then $\sigma(Y)\subset\mathcal{G}\setminus\mathcal{G}^{(0)}$ is a nontrivial bisection and $|C_{g}|<+\infty$ for every $g\in\sigma(Y)$ by construction.
@@
Let $(\mathcal{G},\mu)$ be a discrete Borel groupoid. Write $\Gamma_{x}=\{g\in\mathcal{G}:s(g)=t(g)\}$ for $x\in\mathcal{G}^{(0)}$ and for an element $g\in\Gamma_{x}$ write $C_{g}\subset\Gamma_{x}$ for its conjugacy class. The sets $B=\{x\in\mathcal{G}^{(0)}\mid\Gamma_{x}\text{ is trivial}\}$ and $A=\{x\in\mathcal{G}^{(0)}\mid\Gamma_{x}\text{ is icc}\}$ are Borel and there exists a Borel section $\sigma\colon\mathcal{G}^{(0)}\rightarrow\operatorname{Iso}(\mathcal{G})$ such that $\sigma(x)$ is nontrivial and has finite conjugacy class for every $x\in\mathcal{G}^{(0)}\setminus A$.
@@
The source map $s\colon\operatorname{Iso}(\mathcal{G})\rightarrow\mathcal{G}^{(0)}$ is a surjective and countable-to-one Borel map. So by the Lusin–Novikov theorem there exists a countable collection of Borel subsets $U_{n}\subset\operatorname{Iso}(\mathcal{G})$ such that $\operatorname{Iso}(\mathcal{G})=\bigcup_{n}U_{n}$ and such that $s\colon U_{n}\rightarrow\mathcal{G}^{(0)}$ is a Borel isomorphism for every $n$. Note that $x\in B$ if and only if $x\in U_{n}$ for every $n$, where we identify $x$ and $e_{x}$. Then we have that $B=\bigcap_{n}(\mathcal{G}^{(0)}\cap U_{n})$, which is Borel.
The restriction of $s$ to $\operatorname{Iso}(\mathcal{G})\setminus\mathcal{G}^{(0)}$ is again countable-to-one, so we find countably many Borel section $\tau_{n}\colon\mathcal{G}^{(0)}\setminus B\rightarrow\operatorname{Iso}(\mathcal{G})$ such that $\operatorname{Iso}(\mathcal{G})\setminus\mathcal{G}^{(0)}=\bigcup_{n}\tau_{n}(\mathcal{G}^{(0)})$. For $g\in\Gamma_{x}$ write $C_{g}\subset\Gamma_{x}$ for its conjugacy class. If $\sigma\colon\mathcal{G}^{(0)}\rightarrow\operatorname{Iso}(\mathcal{G})$ is a Borel section, then the subset $\bigsqcup_{x\in\mathcal{G}^{(0)}}C_{\sigma(x)}=\bigcup_{n}U_{n}\sigma(\mathcal{G}^{(0)})U_{n}^{-1}\subset\operatorname{Iso}(\mathcal{G})$ is Borel. Then, since for any countable-to-one Borel map $q$, the assignment $x\mapsto|q^{-1}(x)|$ is Borel, we get that $x\mapsto|C_{\tau_{n}(x)}|$ is Borel for every $n$. Now define Borel maps
$f_{n}\colon\mathcal{G}^{(0)}\setminus B\rightarrow\mathbb{N}\cup\{+\infty\}:\quad f_{n}(x)=\begin{cases}+\infty&\text{ if }|C_{\tau_{n}(x)}|=+\infty,\\ n&\text{ if }|C_{\tau_{n}(x)}|<+\infty,\end{cases}$ (18)
@@
$\sigma(x)=\begin{cases}\tau_{n}(x)&\text{ if }f(x)=n<+\infty,\\ e_{x}&\text{ if }x\in A.\end{cases}$ (19)
Then $\sigma(x)$ is nontrivial and has finite conjugacy class for every $x\in\mathcal{G}^{(0)}\setminus A$. ∎
@@ 6.2 Transformation groupoids
As mentioned in the introduction Vaes provided necessary and sufficient conditions for when the crossed $L^{\infty}(X,\mu)\rtimes G$ by a nonsingular group action $G\curvearrowright(X,\mu)$ is a factor in [Vae20]. In this subsection we deduce this characterization as a corollary of Theorem A, see Corollary 6.4 below. In particular, [Vae20, Corollary] says that if the action $G\curvearrowright(X,\mu)$ is moreover pmp, the crossed product $L^{\infty}(X,\mu)\rtimes G$ is a factor if and only if $G\curvearrowright(X,\mu)$ is ergodic and every $g\in G$ satisfying $\mu(\{x\in X:g\cdot x=x\})>0$ has an infinite conjugacy class in $G$.
Recall that to a nonsingular action $G\curvearrowright(X,\mu)$ of a countable group $G$ on a standard probability space $(X,\mu)$ we associate the transformation groupoid $G\ltimes X$. As a set, we have that $G\ltimes X=G\times X$. The set of composable pairs is given by $(G\ltimes X)^{(2)}=\{((g,h\cdot x),(h,x))\mid g,h\in G,x\in X\}$ and multiplication is defined by $(g,h\cdot x)(h,x)=(gh,x)$. Identifying the unit space $(G\ltimes X)^{(0)}$ with $X$, we get that $s(g,x)=x$ and $t(g,x)=g\cdot x$. There is a canonical isomorphism $L^{\infty}(X,\mu)\rtimes G\cong L(G\ltimes X)$ given by $f\mapsto f$ for $f\in L^{\infty}(X,\mu)$ and $u_{g}\mapsto\lambda_{\{g\}\times X}$ for $g\in G$, where the $u_{g}$ denote the canonical unitaries generating $L(G)$ inside $L^{\infty}(X,\mu)\rtimes G$.
@@
First assume that $\mathcal{G}$ is not icc. Then there exists a nontrivial bisection $B\subset\operatorname{Iso}(\mathcal{G})\setminus\mathcal{G}^{(0)}$ which has $\mu$-finite conjugacy class $\Omega_{B}$. Denote by $\mathcal{P}(G)$ the power set of $G$ and define the function $F\colon X\rightarrow\mathcal{P}(G)$ by
@@
Write $Z=\{x\in X\mid F(x)\neq\emptyset\}$. Then $\mu(Z)>0$ since $\mu(s(B))>0$. Furthermore, since $\Omega_{B}$ is conjugation-invariant, we have that $F(g\cdot x)=gF(x)g^{-1}$ for every $x\in X$ and $g\in G$. In particular, $Z$ is $G$-invariant and we get that $\mu(Z)=1$. Also we have that $F(x)\subset G\setminus\{e\}$ for every $x\in Z$ because $B\cap\mathcal{G}^{(0)}=\emptyset$. Since $\mu_{s}(\Omega_{B})=\int_{X}|F(x)|\mathrm{d}\mu(x)$, we have that $|F(x)|<+\infty$ for a.e. $x\in Z$. There are only countably many nonempty finite subsets $C\subset G\setminus\{e\}$, so we can pick a finite nonempty set $C\subset G\setminus\{e\}$ such that the set $Y=\{x\in X:F(x)=C\}$ has positive measure. Then we have that $\mu(gY\cap Y)>0$ if and only if $gCg^{-1}=C$ if and only if $gY=Y$. So, if we write $H=\{h\in G:hCh^{-1}=C\}$ we get that $G\curvearrowright X$ is induced from $H\curvearrowright Y$. When $g\in C$ and $y\in Y$, we have that $11_{\Omega_{B}}(g,y)=$ 1. This means that $(g,y)\in\Omega_{B}\subset\operatorname{Iso}(\mathcal{G})$, so that $g\cdot y=y$. In particular we have that $F(g\cdot y)=F(y)$ and it follows that $g\in H$. So $C$ is a finite subset of $H\setminus\{e\}$ that is invariant under conjugation by $H$ and all its elements act trivially on $Y$, contradicting statement $1$ of the proposition.
@@
Using that $L(G\ltimes X)\cong L^{\infty}(X,\mu)\rtimes G$ and that $L^{\infty}(X,\mu)^{G}\subset\mathcal{Z}(L^{\infty}(X,\mu)\rtimes G)$ we thus obtain the following as a corollary of Theorem A and Proposition 6.3.
@@
    The crossed product $L^{\infty}(X,\mu)\rtimes G$ is a factor.
@@ 6.3 Full subgroupoids and partial actions
In the previous section we have demonstrated that factoriality of crossed products $L^{\infty}(X,\mu)\rtimes G$ is inherently linked to factoriality of certain subgroupoids of $G\ltimes X$. First we describe subgroupoids which share the same ideal structure as our original groupoid.
@@
As an example, when $\mathcal{G}$ is an equivalence relation a subset $K\subset\mathcal{G}^{(0)}$ is full precisely when it intersects every orbit.
For what follows, we let $\mathcal{R}$ denote the full equivalence relation on the set $\mathbb{N}$ of positive integers. Our goal is to show that subgroupoids induced from full subsets of $\mathcal{G}^{(0)}$ share the same ideal structure as $\mathcal{G}$. What follows is essentially the same proof as in [Car17, Lemmas 2.2, 2.3, 2.4, Proposition 2.5] in the ample groupoid case.
@@
The proof is a back and forth construction using Lemma 6.6, and is identical to the proof in [Car17, Lemma 2.4], so we omit the details here. ∎
@@
$\mathcal{G}\times\mathcal{R}\to\mathcal{G}|_{K}\times\mathcal{R}:g\mapsto W^{-1}gW\;,$
where $W^{-1}gW=w^{-1}gw$ for the unique $w\in W$ with $t(g)=t(w)$, gives us the groupoid isomorphism, with inverse $g\mapsto WgW^{-1}$. ∎
@@
Consider the identity $L(\mathcal{G})\overline{\otimes}B(\ell^{2}(\mathbb{N}))=L(\mathcal{G}\times\mathcal{R})\cong L(\mathcal{G}|_{K}\times\mathcal{R})=L(\mathcal{G}|_{K})\overline{\otimes}B(\ell^{2}(\mathbb{N}))$. Since the center of the tensor product of von Neumann algebras is the tensor product of their centers, the result now follows from Theorem A. ∎
@@
If $G\curvearrowright X$ is induced from $H\curvearrowright Y$, then $Y$ is full by definition. Conversely, note that by fullness of $Y$, it follows that $\bigcup_{g\in G}gY=t((G\ltimes X)Y)=X$. To show the sets $gY$ are mutually disjoint, let $g\not\in H$. If there is some $y\in gY\cap Y$, then $(g,y)\in(G\ltimes X)|_{Y}=H\ltimes Y$. Thus, $g\in H$, reaching a contradiction. The measurable setting is the same, with equality replaced with equality up to a negligable subset. ∎
Note that subgroupoids of $G\ltimes X$ are not necessarily crossed products. Indeed, even induced subgroupoids of $G\ltimes X$ are not necessarily a crossed product.
@@
To get around this issue, one can consider at partial actions instead. The following is a translation of the partial dynamical system construction in the topological case to the Borel setting. See the book [Exe17] and the references therein for more details.
@@
Set $Y_{g}=\{x\in X_{g}\;|\;x\in Y,\sigma_{g^{-1}}(x)\in Y\}$. For any $x\in Y_{g}$, $\sigma_{g^{-1}}(x)\in Y_{g^{-1}}$. It follows that $\sigma_{g}\colon Y_{g^{-1}}\rightarrow Y_{g}$ is a bijection with inverse $\sigma_{g^{-1}}$. The only non-trivial fact to verify is that $(\sigma_{g}\circ\sigma_{h})(x)\in Y_{gh}$ if $x\in\sigma_{h}^{-1}(Y_{h}\cap Y_{g^{-1}})$. Since $(\sigma_{g}\circ\sigma_{h})(x)=\sigma_{gh}(x)\in Y\cap X_{gh}$, it is enough to verify that $x\in Y_{(gh)^{-1}}$. This follows from the fact that $x,\sigma_{gh}(x)\in Y$. Finally, to see that $(G\ltimes X)|_{Y}=G\ltimes Y$, observe that $\{g\}\times Y_{g^{-1}}$ form a basis for both groupoids. ∎
Indeed, partial actions always come from the restriction of a global action. The proof is a modification of [Aba03, Theorem 2.5] in the topological case.
@@
Let $G\curvearrowright_{\sigma}Y$ be a partial action on a standard Borel space $Y$. Then there is an embedding $Y\hookrightarrow X$ into a standard Borel space $X$ and a global action $G\curvearrowright_{\theta}X$ such that $G\ltimes Y=(G\ltimes X)|_{Y}$. Furthermore, $X$ can be chosen such that $Y$ is $G\ltimes X$-full.
@@
Next, we show by induction that $[(g,x)]\cap A$ is nonempty for every $(g,x)\in G\times Y$. Take an arbitrary element $(g_{i},x)\in G\times Y$. If $i=0$ we are done. Now assume that $i>0$. If $x\notin\cup_{k=0}^{i-1}Y_{g_{i}^{-1}g_{k}}$, then we have that $(g_{i},x)\in A$ by definition. If $x\in\cup_{k=0}^{i-1}Y_{g_{i}^{-1}g_{k}}$, there exists a $k<i$ such that $x\in Y_{g_{i}^{-1}g_{k}}$ and we get that $(g_{i},x)R(g_{k},\sigma_{g_{i}^{-1}g_{k}}(x))$. By the induction hypothesis we get that $[(g_{i},x)]\cap A=[(g_{k},\sigma_{g_{i}^{-1}g_{k}}(x))]\cap A\neq\emptyset$.